# Graph Quantum Magic Squares and Free Spectrahedra

### v2

@@ 4 Future Directions
4.2 (2) Arveson extreme points of $\mathcal{M}^{(\Gamma)}$
For the non-graph case, [DDN20, EH19] show that
@@
In the graph setting, our Corollary 3.11 establishes the analogous extremality statement:
@@
What remains open is the following question:
- Can one characterise the full family of Arveson extreme points?
Understanding the full extremal structure of $\mathcal{M}^{(\Gamma)}$ is central for determining its “vertices” in the matrix convex sense, and for comparing the graph case with the non-graph setting of [DDN20].
4.3 (3) Connections with quantum information theory
Quantum magic squares are deeply connected with non-local games and projective/positive-operator valued measurements (PVM vs. POVM). In particular, the failure of the Birkhoff–von Neumann property has been conjecturally linked with PVM-POVM separation phenomena in graph automorphism games.
Possible future directions include:
- establishing a precise correspondence between GQMS and the strategy sets of graph non-local games;
- identifying graphs for which $\mathcal{M}^{(\Gamma)}$ witnesses PVM $\neq$ POVM behaviour.
These connections would further place graph quantum magic squares within the framework of quantum information theory.
@@ A.1 Dual SDP and Explicit Certificate
In this appendix we briefly explain the semidefinite program used to produce the dual certificate $Y$ in the proof of Theorem 3.3.
@@
is the Hermitian subspace appearing in Proposition 2.5. If $B\in\mathrm{mconv}(\mathcal{P}^{(4)})_{2}$, then there exists $X\in\mathcal{S}$ such that $M_{B}+X\succeq 0$.
@@
Thus, the condition $M_{B}+X\succeq 0$ can be reformulated as the existence of coefficients $\xi_{a,b,j}$ such that
@@
This provides the dual certificate used in Theorem 3.3.
@@ A.1.1 Explicit basis for $\mathcal{Z}^{(4)}_{e}$
For completeness, we record an explicit Hermitian basis of the subspace
@@
which has dimension $5$. A convenient choice is given by the following matrices:
@@
Writing $Y_{i,j,k}:=z_{i}\otimes z_{j}\otimes s_{k}$, an (explicit) Hermitian spanning family of $\mathcal{S}=(\mathcal{Z}^{(4)}_{e}\otimes\mathcal{Z}^{(4)}_{e}\otimes\operatorname{Mat}_{2}\!\left(\mathbb{C}\right))_{\mathrm{her}}$ is obtained via
@@
and arbitrary real linear combinations of the $C_{i,j,k}$ and $D_{i,j,k}$. (Any other basis of $\mathcal{Z}^{(4)}_{e}$ works.)
We now compute the dimension of the commutant of the adjacency matrix for cyclic graphs $C_{n}$, as this determines the number of independent Hermitian parameters appearing in the affine parametrization of the $C_{n}$-quantum magic squares.
@@ A.2 Explicit computation of the commutant dimension for cycle graphs
In this appendix, we compute this dimension explicitly for the family of cycle graphs $C_{n}$, which corresponds to the case of the counterexample discussed in Section 3.2.
@@ A.2.1 Cycle graphs $C_{n}$
We now compute $d_{n}=\dim\{X\in\operatorname{Mat}_{n}(\mathbb{C})\mid(A_{C_{n}}\otimes I_{s})X=X(A_{C_{n}}\otimes I_{s})\}$ for the cycle graph $C_{n}$, which is connected and $2$-regular.
@@
Let $A_{C_{n}}\in\operatorname{Mat}_{n}(\mathbb{R})$ be the adjacency matrix of the cycle $C_{n}$. Set
@@
Since $A_{C_{n}}$ is real symmetric, it is diagonalizable; the dimension of its commutant in $\operatorname{Mat}_{n}(\mathbb{C})$ equals $\sum_{\lambda}m_{\lambda}^{2}$, where $m_{\lambda}$ is the multiplicity of the eigenvalue $\lambda$.
@@
For $n$ odd, $\lambda_{0}=2$ has multiplicity $1$, and the remaining indices occur in pairs (k,nk)(k,n-k) with $1\leq k\leq(n-1)/2$, each contributing an eigenspace of multiplicity $2$. Thus
@@
For $n$ even, $\lambda_{0}=2$ and $\lambda_{n/2}=-2$ are simple; the remaining indices pair as (k,nk)(k,n-k) with $1\leq k\leq n/2-1$, each with multiplicity $2$. Hence
@@
For a QMS commuting with $A_{C_{n}}$ (connected case), the number of independent Hermitian block variables equals $d_{n}-1$, i.e.,
@@
The subtraction of $1$ accounts for the single component ($N=1$) and the resulting componentwise bistochastic constraint (row/column sums equal to $I_{s}$). In the even case, the pairing of indices (k,nk)(k,n-k) is already reflected in the multiplicity pattern that yields $d_{n}$.

### v3

@@ 4.1 Beyond $C_{4}$: the (Quantum) Birkhoff–von Neumann problem for graphs
4.2 Arveson extreme points of $\mathcal{M}^{(\Gamma)}$
For the QMS case, [DDN20, EH19] show that
@@ 4.2 Arveson extreme points of $\mathcal{M}^{(\Gamma)}$
For the GQMS case, Corollary 3.12 establishes the analogous statement:
@@
Then we can ask whether we can characterise the full family of Arveson extreme points.
To identify the “vertices” of $\mathcal{M}^{(\Gamma)}$, we need a better understanding of its extremal structure, and this also helps us compare the graph and non-graph settings in [DDN20].
4.3 Connections with quantum information theory
Quantum magic squares can be formulated in terms of positive operator-valued measures (POVMs) and thus connect to quantum information theory, for example via non-local games. Since quantum permutation matrices correspond to projection-valued measurement (PVM), it is natural to ask whether the failure of a Birkhoff-von Neumann theorem type for quantum magic squares can be interpreted as a separation between POVM and PVM.
Hence, a natural direction for future work is to investigate this perspective for Graph Quantum Magic Squares (GQMS) and to identify graphs $\Gamma$ for which $\mathcal{M}^{(\Gamma)}$ exhibits genuinely different behaviors in the PVM vs POVM models. Clarifying these points would further connect GQMS to quantum information theory.
@@ A.1 Dual SDP and Explicit Certificate
In this appendix we briefly explain the semidefinite program used in the proof of Theorem 3.3.
@@
is the Hermitian subspace of Proposition 2.5. If $B\in\mathrm{mconv}(\mathcal{P}^{(4)})_{2}$, then there exists $X\in\mathcal{S}$ such that $M_{B}+X\succeq 0$.
@@
Thus, the condition $M_{B}+X\succeq 0$ can be reformulated in terms of coefficients $\xi_{a,b,j}$ such that
@@
This provides the construction of the matrix $Y$ used in Theorem 3.3.
@@ A.1.1 Explicit basis for $\mathcal{Z}^{(4)}_{e}$
For completeness, we give an explicit Hermitian basis of the subspace
@@
which has dimension $5$. We chose the following matrices:
@@
Writing $Y_{i,j,k}:=z_{i}\otimes z_{j}\otimes s_{k}$, we can describe an explicit Hermitian basis for
$\mathcal{S}=\big(\mathcal{Z}^{(4)}_{e}\otimes\mathcal{Z}^{(4)}_{e}\otimes\operatorname{Mat}_{2}\!\left(\mathbb{C}\right)\big)_{\mathrm{her}}$
by setting
@@
so that $\mathcal{S}$ is spanned by arbitrary real linear combinations of $C_{i,j,k}$ and $D_{i,j,k}$. (Any other basis of $\mathcal{Z}^{(4)}_{e}$ would work as well.)
As an example, we now compute the dimension of the commutant of the adjacency matrix for cyclic graphs $C_{n}$.
@@ A.2 Explicit computation of the commutant dimension for cycle graphs
Here, we compute this dimension explicitly for the cycle graphs $C_{n}$, which corresponds to the case of the counterexample discussed in Section 3.2.
@@ A.2.1 Cycle graphs $C_{n}$
Let $d_{n}=\dim\{X\in\operatorname{Mat}_{n}(\mathbb{C})\mid(A_{C_{n}}\otimes I_{s})X=X(A_{C_{n}}\otimes I_{s})\}$ for $C_{n}$, and recall that $C_{n}$ is connected and $2$-regular.
@@
Let $A_{C_{n}}\in\operatorname{Mat}_{n}(\mathbb{R})$ be the adjacency matrix of $C_{n}$ and set
@@
Since $A_{C_{n}}$ is real symmetric, it is diagonalizable; the dimension of its commutant in $\operatorname{Mat}_{n}(\mathbb{C})$ is $\sum_{\lambda}m_{\lambda}^{2}$, where $m_{\lambda}$ is the multiplicity of the eigenvalue $\lambda$.
@@
For $n$ odd, $\lambda_{0}=2$ has multiplicity $1$. The other indices appear in pairs (k,nk)(k,n-k) with $1\leq k\leq(n-1)/2$, each of them with multiplicity $2$. Thus
@@
For $n$ even, $\lambda_{0}=2$ and $\lambda_{n/2}=-2$ have multiplicity $1$. The remaining indices pair as (k,nk)(k,n-k) with $1\leq k\leq n/2-1$, each with multiplicity $2$. Hence
@@
For a QMS commuting with $A_{C_{n}}$, the number of independent Hermitian block variables is $d_{n}-1$, i.e.,
@@
The subtraction of $1$ takes into account the single component ($N=1$) and the resulting componentwise doubly stochastic constraint (row/column sums equal to $I_{s}$). In the even case, the index pairing (k,nk)(k,n-k) is already reflected in the multiplicity pattern that produces $d_{n}$.